The triangle is isosceles and the perimeter equals 54 cm. The height is 3*sqrt(21). Find the lengths of the sides and the area of the triangle. Express your answers in simplest radical form.
Answer by TimothyLamb(4379)

definition of isosceles triangle: a triangle with two equal sides
---
the problem doesn't say which height is being talked about, so let's assume it's the height from the non-equal side (the base b), to the point where the two equal sides meet:
---
perimeter = 54 = 2s + b
2s = 54 - b
s = (54 - b)/2
---
length of the equal sides:
---
s = sqrt( hh + bb/4 )
(54 - b)/2 = sqrt( hh + bb/4 )
(54 - b)(54 - b)/4 = hh + bb/4
54*54 - 108b + bb = 4hh + bb
2916 - 108b = 4hh
108b = 2916 - 4hh
b = (2916 - 4hh)/108
b = (2916 - 4* 3*sqrt(21) * 3*sqrt(21) )/108
b = (2916 - 4*3*3*sqrt(21)*sqrt(21) )/108
b = (2916 - 4*3*3*21 )/108
b = 20 cm
s = (54 - b)/2
s = (54 - 20)/2
s = 17 cm
---
area:
---
a = bh/2
a = 20*3*sqrt(21)/2
a = 30*sqrt(21) sq.cm
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---